### v1

@@ I Introduction
represents a seminal result in our understanding of the origin of $CP$ violation within the Standard Model. Since CPV in both mesons and baryons arises from the CKM mechanism [19, 20] that governs quark transitions, a crucial test of the theory comes from flavor symmetries, particularly given the challenges in making precise theoretical predictions for $CP$ violation in multibody baryon decays. Several approaches have been proposed to study $CP$ violation in multibody bottom baryon decays [21, 22, 23], including a method to extract the $CP$-violating weak phase in charmless $b$-baryon decays [24, 25].
@@ II Application of $U$-spin to decay amplitudes
×$\times C^{U_{3}^{M}\,U_{3}^{B_{1}}\,U_{3}^{F}}_{U^{M}\,U^{B_{1}}\,U^{F}}C^{U_{3}^{F}\,(-U_{3}^{B_{b}})\,U_{3}^{\mathcal{H}}}_{U^{F}\,U^{B_{b}}\,U^{\mathcal{H}}}A^{U^{M},U^{F},\pm}(U^{\mathcal{H}})$
@@ III Amplitude Relations
$\Lambda^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$ $\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$
$\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}$
@@
where $i$, $j$, and $a,\,e$ $b,\,f$, $c,\,g$, $d,\,h$ indices correspond to the various baryon and mesons in the initial and final states and the negative sign accounts for the relative sign between the Jarlskog factor in $\Delta|S|=1$ and $\Delta S=0$ decay modes. These $A_{CP}$-relations can be read off from Table (2). In particular, measurements for the following branching ratios of $\Lambda^{0}_{b}$-baryon [44],
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})=(5.0\pm 0.5)\times 10^{-5}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}K^{+})=(4.0\pm 0.6)\times 10^{-6}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})=(1.25\pm 0.13)\times 10^{-5}$ (16)
$Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})=(2.08\pm 0.21)\times 10^{-5}$
can be used to express the $CP$-asymmetries in the decay modes $\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$, $\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+}$, $\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ and $\Xi^{0}_{b}\to\Sigma^{+}K^{-}K^{-}K^{+}$ respectively.
@@
We now discuss the effect of first order breaking of $U$-spin originating form unequal masses of the $d$ an $s$ quark. $U$-spin breaking is introduced by a triplet under $U$-spin ($U_{\text{brk}}=1,U^{3}_{\text{brk}}=0$) and by taking the tensor product with the unbroken Hamiltonian that transforms as a doublet. The effect of this $U$-spin breaking is a correction to the unbroken Hamiltonian given by,
@@
$A_{\text{brk}}^{\frac{1}{2},0,\pm}(\frac{1}{2})=\frac{\epsilon_{1}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},0,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{2}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{3}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{3}{2})$ (18)
$A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{4}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{5}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{3}{2})$
$A_{\text{brk}}^{\frac{3}{2},2,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{6}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},2,\pm}(\frac{3}{2})$
The combined result of leading order unbroken Hamiltonian and first order $U$-spin breaking contribution to the decay amplitudes is given in Appendix (B). The first order $U$-spin breaking corrections have the same structures but opposite signs for the decay modes related by $U$-spin as found in [29]. This increases the rank of the coefficient matrices to 12 and the leading order relations between $U$-spin conjugate amplitude pairs no longer hold. Naturally, the leading order $CP$-asymmetry relations evident from Table (2) also breaks down once $U$-spin breaking effects are considered. Any reduction among the set of $U$-spin subamplitudes would require knowledge of their relative magnitudes, which lies beyond the scope of the present work. We note that if a single $U$-spin subamplitude dominates, the framework becomes significantly more predictive.
@@ IV Conclusion
We have investigated charmless bottom baryon decays into fully charged four-body final states within the realm of $U$-spin symmetry. Bearing in mind that significant $U$-spin breaking is observed in $B$-meson decays, we set up a framework to analyze bottom baryon decays where $U$-spin symmetry can be arbitrarily broken. We obtained amplitude relations between decay modes and based on those relations predicted $CP$ asymmetry relations between a number of decay modes. While $U$-spin breaking effect spoils the relations between observables, the predictions offer a roadmap for probing $CP$ violation in bottom baryon decays and provide a systematic approach to test the validity of $U$-spin symmetry in future experimental measurements.
@@
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&-\frac{1}{2}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&\frac{1}{3}&-\frac{1}{3}&\frac{1}{3\sqrt{2}}&-\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&\frac{2}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{6}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{5}{2})\end{pmatrix}$ (33)

### v2

@@ I Introduction
represents a seminal result in our understanding of the origin of $CP$ violation within the Standard Model. Since CPV in both mesons and baryons arises from the CKM mechanism [19, 20] that governs quark transitions, a crucial test of the theory comes from flavor symmetries, particularly given the challenges in making precise theoretical predictions for $CP$ violation in multibody baryon decays. Several approaches have been proposed to study $CP$ violation in multibody bottom baryon decays [21, 22, 23], including a method to extract the $CP$-violating weak phase in charmless bottom baryon decays [24, 25].
@@ II Application of $U$-spin to decay amplitudes
$\times C^{U_{3}^{M}\,U_{3}^{B_{1}}\,U_{3}^{F}}_{U^{M}\,U^{B_{1}}\,U^{F}}C^{U_{3}^{F}\,(-U_{3}^{B_{b}})\,U_{3}^{\mathcal{H}}}_{U^{F}\,U^{B_{b}}\,U^{\mathcal{H}}}A^{U^{M},U^{F},\pm}(U^{\mathcal{H}})$
@@ III Amplitude Relations
$\Lambda^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$ $\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ $\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}$
@@
where $i$, $j$, and a,ea,\,e b,fb,\,f, c,gc,\,g, d,hd,\,h indices correspond to the various baryon and mesons in the initial and final states and the negative sign accounts for the relative sign between the Jarlskog factor in $\Delta|S|=1$ and $\Delta S=0$ decay modes. These $A_{CP}$-relations can be read off from Table (2) by using Eq (15). As reported by LHCb, the $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ decay as well as other four-body bottom baryon decays occur primarily through hadronic resonances that individually decay into two or three final-state particles. The contribution from multiple resonances results in substantial variation in associated strong phases and relative strengths of tree ($\mathcal{A}_{T}$) and penguin amplitudes ($\mathcal{A}_{P}$) leading to differences in the $CP$-asymmetry across the final-state phase space of the bottom baryon decay. Therefore, it is reasonable to investigate local $CP$-asymmetries near resonant regions in such decays, as will be explored in the following. If all the final-state mesons originate from a system with $U$-spin = 1/2, as observed in one of the four phase-space regions of the decay $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ [1], the $A^{\frac{3}{2},1,\pm}(\frac{1}{2})$ $U$-spin subamplitudes can be dropped. This simplification allows us to describe decays in Table (2) in terms of 4 $U$-spin subamplitudes, two for each class of decays with a positively or negatively charged baryon in the final state. This leads to additional $CP$-asymmetry relations that are given by,
$A_{CP}(\Lambda^{0}_{b}\to p^{+}(K^{-}K^{-}K^{+}))$ $\simeq A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}K^{-}\pi^{+})),$
$=-A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(\pi^{-}\pi^{-}\pi^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})}{Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})}$
$=-\frac{1}{2}A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(\pi^{-}K^{-}K^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}K^{+})}{Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})}$ (16a)
$A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}\pi^{-}\pi^{+}))$ $\simeq A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}K^{-}K^{+}))$
$=-A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(K^{-}K^{-}K^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})}{Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})}$
$=-\frac{1}{2}A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(\pi^{-}K^{-}\pi^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})}{Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})}$ (16b)
$A_{CP}(\Lambda^{0}_{b}\to\Sigma^{+}(\pi^{-}\pi^{-}\pi^{+}))$ $\simeq A_{CP}(\Lambda^{0}_{b}\to\Sigma^{+}(\pi^{-}K^{-}K^{+}))$
$=-A_{CP}(\Xi_{b}^{0}\to p^{+}(K^{-}K^{-}K^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})}{Br(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})}$
$=-\frac{1}{2}A_{CP}(\Xi_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to p^{+}\pi^{-}K^{-}\pi^{+})}{Br(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})}$ (16c)
Measurements of the following branching ratios of $\Lambda^{0}_{b}$-baryon and $\Xi^{0}_{b}$-baryon are available from LHCb [44],
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})=(5.1\pm 0.5)\times 10^{-5}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}K^{+})=(4.1\pm 0.6)\times 10^{-6}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})=(1.27\pm 0.13)\times 10^{-5}$ (17)
$Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})=(2.12\pm 0.21)\times 10^{-5}$
$Br(\Xi^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\times\frac{F(b\to\Xi_{b}^{0})}{F(b\to\Lambda^{0}_{b})}=(1.9\pm 0.4)\times 10^{-6}$
$Br(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}\pi^{+})\times\frac{F(b\to\Xi_{b}^{0})}{F(b\to\Lambda^{0}_{b})}=(1.73\pm 0.31)\times 10^{-6}$
$Br(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\times\frac{F(b\to\Xi_{b}^{0})}{F(b\to\Lambda^{0}_{b})}=(1.8\pm 1.0)\times 10^{-7}$
where $F(b\to B_{b})$ denotes the fragmentation fraction of $b$-quarks into a $B_{b}$-baryon. From the available local $CP$-asymmetry measurements in bottom baryon decays [13, 1], we identify that $CP$-asymmetry in $\Lambda_{b}^{0}\to p^{+}K_{1}^{-}(1410)$ quasi-two body decay [13],
$A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+}))=(4.7\pm 3.5\pm 0.8)\%,$ (18)
can be translated into local $CP$-asymmetry in $\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}\pi^{+}$ transition since $K_{1}^{-}(1410)$ decays strongly to $\pi^{+}K^{-}\pi^{-}$ [44]. Moreover, $K_{1}^{-}(1410)\to K^{+}K^{-}K^{-}$ strong decay is also possible [45] and therefore based on Eq (16a) we expect that local $CP$-asymmetry in $\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ transition is,
$A_{CP}(\Lambda_{b}^{0}\to p^{+}(K^{-}K^{-}K^{+})_{K_{1}})=A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+})_{K_{1}})\simeq(4.7\pm 3.6)\%$ (19)
We observe that the predicted value of $A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+})_{K_{1}})$ has the same sign as the recently measured local $A_{CP}$ [1]
$A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+}))|_{m_{\pi^{+}K^{-}\pi^{-}}<\,2.0\,\text{GeV}}=(2.0\pm 1.2\pm 0.3)\%,$ (20)
although the central values differ by roughly a factor of two. Similarly, we focus on $CP$-violation in the quasi-two body $\Lambda^{0}_{b}\to p^{+}a_{1}^{-}(1260)$ decay where we note that $a_{1}^{-}(1260)$ subsequently can decay strongly to either $\pi^{+}\pi^{-}\pi^{-}$ or $K^{+}\pi^{-}K^{-}$ [44]. We use the measured $CP$-asymmetry in $\Lambda^{0}_{b}\to p^{+}a_{1}^{-}(1260)$-decay [13],
$A_{CP}(\Lambda^{0}_{b}\to p^{+}a_{1}^{-}(1260))=(-1.5\pm 4.2\pm 0.6)\%$ (21)
and with the help of Eq (16b) also predict local $CP$-asymmetry in $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}K^{+}$ and $\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+}$,
$A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}\pi^{-}\pi^{+})_{a_{1}})=A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}K^{-}K^{+})_{a_{1}})\simeq(-1.5\pm 4.2)\%.$ (22)
@@
Once the branching fractions and local $CP$ asymmetries of additional $\Lambda^{0}_{b}$ and $\Xi^{0}_{b}$ decay modes, such as $\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ and $\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}K^{-}K^{+}$ become available, they can be used to express the local-$CP$ asymmetries in related $\Xi^{0}_{b}$ decay modes: $\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ and $\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ with the help of Eq (16c). The $CP$-asymmetry relations inferred from Table (2) can also be tested by experiments in near future. We now discuss the effect of first order breaking of $U$-spin originating form unequal masses of the $d$ an $s$ quark. $U$-spin breaking is introduced by a triplet under $U$-spin ($U_{\text{brk}}=1,U^{3}_{\text{brk}}=0$) and by taking the tensor product with the unbroken Hamiltonian that transforms as a doublet. The effect of this $U$-spin breaking is a correction to the unbroken Hamiltonian given by,
@@
$A_{\text{brk}}^{\frac{1}{2},0,\pm}(\frac{1}{2})=\frac{\epsilon_{1}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},0,\pm}(\frac{1}{2}),\qquad A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{2}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{3}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{3}{2}),\qquad A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{4}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{1}{2}),$ (24)
$A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{5}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{3}{2}),\qquad A_{\text{brk}}^{\frac{3}{2},2,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{6}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},2,\pm}(\frac{3}{2})$
The combined result of leading order unbroken Hamiltonian and first order $U$-spin breaking contribution to the decay amplitudes is given in Appendix (B). The first order $U$-spin breaking corrections have the same structures but opposite signs for the decay modes related by $U$-spin as found in [29]. This increases the rank of the coefficient matrices to 12 and the leading order relations between $U$-spin conjugate amplitude pairs no longer hold. Naturally, the leading order $CP$-asymmetry relations evident from Table (2) and Eq (16) also break down once $U$-spin breaking effects are considered. Any reduction among the set of $U$-spin subamplitudes would require knowledge of their relative magnitudes, which lies beyond the scope of the present work.
@@ IV Conclusion
We have investigated charmless bottom baryon decays into fully charged four-body final states within the realm of $U$-spin symmetry. Bearing in mind that significant $U$-spin breaking is observed in $B$-meson decays, we set up a framework to analyze bottom baryon decays where $U$-spin symmetry can be arbitrarily broken. We obtained amplitude relations between decay modes and based on those relations predicted $CP$ asymmetry relations between a number of decay modes. By making the additional assumption that all final-state mesons emerge from a $U$-spin doublet, we demonstrated that this extended framework allows for numerical predictions of localized $CP$ asymmetries in specific regions of phase space, which hold to a good accuracy in practice. The framework also predicts the relative signs of the $CP$ asymmetries, an important feature that can be obscured in global $CP$ asymmetry measurements due to averaging over the entire phase space of the decay. While $U$-spin breaking effect spoils the relations between observables, the predictions offer a roadmap for probing $CP$ violation in bottom baryon decays and provide a systematic approach to test the validity of $U$-spin symmetry in future experimental measurements.
@@
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&0&0&-\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ 0&\frac{\sqrt{2}}{3}&-\frac{1}{3}&\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&-\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&-\frac{2}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{5}{2})\end{pmatrix}$ (39)